harmony6:

(05.02)The area of the parallelogram below is ____ square meters.

Shadow:

We need data on the parallelogram

harmony6:

(05.02)The area of the parallelogram below is ____ square meters. A parallelogram with height labeled with 7 meters. The top horizontal side is labeled 6 meters. The base of the left triangle formed by the height is 2 meters.

Shadow:

Is there an image associated with this problem?

harmony6:

yes but I can't copy it. I gave the measurements.

harmony6:

1 attachment
Shadow:

There it is!

Shadow:

You can find the area of a parallelogram by multiplying the base by the height. \[A_{p} = bh\]

Shadow:

Area of a parallelogram equals base times height

harmony6:

so I multiply 2x7? What happens to the 6m

Shadow:

You could do 6 x 7 = 42 Or (2 x 7)0.5 = 7 x 2 = 14 + 4 x 7 = 28 28 + 14 = 42

Shadow:

They give you the 2 in case you want to calculate the area of the two triangles adjacent to the 'rectangle' then find the area of the angle, as opposed to just straight calculating the area of the parallelogram.

Shadow:

See the triangles?

Shadow:



harmony6:

Yes and thank you so much for helping me understand how to calculate.

Shadow:



Shadow:

No problem